OpenStudy (anonymous):

Find the sixth term of an arithmetic sequence with t1 = 2 and tn = tn-1 + 4

OpenStudy (amoodarya):

tn-tn-1=4 so t1=2 t2=t1+4=2+4=6 t3=t1+2d .. tn=t1+(n-1)d=2+(n-1)*4 so put n=6